The Good Seed Flower Shop sells its own brand of fertilizer. It mixes one brand of fertilizer that contains 7% nitrogen with another than contains 8% nitrogen. The resulting mixture contains 7.7% nitrogen. How many pounds of the 7% fertilizer must be used to make 400 pounds of the 7.7% mixture?
Answer by checkley77(12844)

.07x+.08(400-x)=.077*400
.07x+32-.08x=30.8
-.01x=30.8-32
-.01x=-1.2
x=-1.2/-.01
x=120 pds. of 7% nitrogen is used.
400-120=380 pds. of 8% is used.
proof:
.07*120+.08*280=30.8
8.4+22.4=30.8
30.8=30.8